what should be the maximum number of electrons in the possible sub shells for n+l = 42

Grade:12th pass

3 Answers

Dhaval
6 years ago
As we know that, n+1=42 is the electrons possible here.. So... When we solve it this way, n=42-1=41We see that the maximum number of electrons in a sub shell.. Are this!!
meghna
6 years ago
no of e =2n2.therefore …........n+1=42 ...therfore n=41..so here n2=412=1681.2*1681=3362.and im not absolutely sure of this answer..i just remember equation like  this..
 
 
ashraful alom
6 years ago
as you know maximum electrons any subshells 2n2 , where , n=2 , 2.4=8 , so the answer is 8